OpenStudy (anonymous):

if log 2 = 0.301 and log 3=0.477 what is the approximate value of log 72?

OpenStudy (anonymous):

log 72=log 36*2=log 36+ log 2 log 36=log 6^2=2log 6= 2(log2+log 3) so log 72= 3log 2 + 2log3= 3*0.301+2*0.477

OpenStudy (campbell_st):

log 72 = log(9 x 8) = log(3^2 x 2^3) using log laws = 2 x log(3) + 3 x log(2)

OpenStudy (accessdenied):

the fact that there was this exact log(72) i left off with in the first problem as a question two years ago is just too amusing to me but hopefully this relic provides some valuable insight.